# Bounding User Contributions for User-Level Differentially Private Mean Estimation

### v2

@@ III-A On Clipping Strategies
We work with estimators $\overline{f}=\overline{f}_{\left\{a_{j}^{(\ell)},b_{j}^{(\ell)}\right\}}$ of $f$ obtained by bounding user contributions as follows: for each $\ell\in[L]$ and $j\in[m_{\ell}]$, we let $\overline{\mathbf{x}}_{j}^{(\ell)}:={\mathsf{A}_{a_{j}^{(\ell)},b_{j}^{(\ell)}}}\left(\mathbf{x}_{j}^{(\ell)}\right)$, for reals $0\leq a_{j}^{(\ell)}\leq b_{j}^{(\ell)}\leq U$. In words, $\overline{\mathbf{x}}_{j}^{(\ell)}$ is a projection of $\mathbf{x}_{j}^{(\ell)}$ onto the set $\mathsf{A}_{a_{j}^{(\ell)},b_{j}^{(\ell)}}$, which, intuitively, reduces the range of values that $\mathbf{x}_{j}^{(\ell)}$ can take, and hence its sensitivity too. Figure 1 shows a pictorial depiction of the annulus $\mathsf{A}_{a_{j}^{(\ell)},b_{j}^{(\ell)}}$ when $d=2$ and examples of projections onto the annulus. We then set $\overline{f}=\overline{f}(\mathcal{D}):=\frac{1}{\sum_{\ell^{\prime}=1}^{L}m_{\ell^{\prime}}}\cdot\sum_{\ell=1}^{L}\sum_{j=1}^{m_{\ell}}\mathbf{\overline{x}}_{j}^{(\ell)}.$
@@ III-B On Worst-Case Errors of Clipping Strategies
For a sample $\mathbf{x}_{j}^{(\ell)}\in\mathsf{A}_{a_{j}^{(\ell)},b_{j}^{(\ell)}}$, it is clear that $\mathsf{A}_{a_{j}^{(\ell)},b_{j}^{(\ell)}}(\mathbf{x}_{j}^{(\ell)})=\mathbf{x}_{j}^{(\ell)}$. Now, suppose that $\mathbf{x}_{j}^{(\ell)}\in\Delta_{U}\setminus\Delta_{b_{j}^{(\ell)}}$. Following Lemma III.1, if $\mathbf{y}=\mathsf{A}_{a_{j}^{(\ell)},b_{j}^{(\ell)}}(\mathbf{a})$, we must have $\lVert\mathbf{y}-\mathbf{x}_{j}^{(\ell)}\rVert_{1}=\sum_{i\leq d}((x_{j}^{(\ell)})_{i}-y_{i})=\lVert\mathbf{x}_{j}^{(\ell)}\rVert_{1}-b_{j}^{(\ell)}.$ Hence, the worst-case clipping error for sample $\mathbf{x}_{j}^{(\ell)}$ is $\max\ \lVert\mathsf{A}_{a_{j}^{(\ell)},b_{j}^{(\ell)}}(\mathbf{x}_{j}^{(\ell)})-\mathbf{x}_{j}^{(\ell)}\rVert_{1}=U-b_{j}^{(\ell)}$. By symmetric arguments, one can show that if $\mathbf{x}_{j}^{(\ell)}\in\Delta_{a_{j}^{(\ell)}}\setminus\delta_{a_{j}^{(\ell)}}$, we must have $\max\ \lVert\mathsf{A}_{a_{j}^{(\ell)},b_{j}^{(\ell)}}(\mathbf{x}_{j}^{(\ell)})-\mathbf{x}_{j}^{(\ell)}\rVert_{1}=a_{j}^{(\ell)}$. Thus, overall, we obtain that
@@
Via arguments entirely analogous to that in the proof of Lemma III.2, the user-level sensitivity of $\overline{f}$ is $\Delta_{\overline{f}}=\frac{\max_{\ell\leq L}\sum_{j\leq m_{\ell}}\left(b_{j}^{(\ell)}-a_{j}^{(\ell)}\right)}{\sum_{\ell^{\prime}\leq L}m_{\ell^{\prime}}},$ since in the worst-case, all samples of a user $\ell$ are changed each from $(a_{j}^{(\ell)},0,\ldots,0)$ to $(b_{j}^{(\ell)},0,\ldots,0)$. Using $\mathbb{E}[\lVert\mathbf{Z}\rVert_{1}]=d\mathbb{E}[|Z_{1}|]=d\Delta_{\overline{f}}/\varepsilon$ gives us the required result. ∎
@@
The optimization problem in (5) is a linear programming problem. By standard arguments via the necessity of the KKT conditions [18, Sec. 5.5.3], there must exist reals $\lambda_{\tau}\geq 0$ and $\lambda_{\ell},\ \mu_{\ell}\geq 0$, for each $\ell\in[L]$ (or Lagrange multipliers), such that the function ($\mathcal{L}(\{S^{(\ell)}\},\ \lambda_{\tau},\{\lambda_{\ell},\mu_{\ell}\}):=\sum_{\ell\leq L}S^{(\ell)}+\frac{d\tau}{\varepsilon}-\lambda_{\tau}\tau\ +\\ \sum_{\ell}\lambda_{\ell}\cdot\left(Um_{\ell}-2S^{(\ell)}-\tau\right)-\sum_{\ell}\mu_{\ell}S^{(\ell)}$ obeys the following properties.
@@
  • Complementary slackness: We have that $\lambda_{\tau}\tau=0,\ \lambda_{\ell}\cdot\left(Um_{\ell}-2S^{(\ell)}-\tau\right)=0,\ \text{and}\ \mu_{\ell}S^{(\ell)}=0,$ for all $\ell\in[L]$.
@@
Observe that from Theorem III.1 and (4), the optimal worst-case error is $E^{\text{OPT}}(\varepsilon)=\frac{1}{\sum_{\ell^{\prime}\leq L}m_{\ell^{\prime}}}\cdot\left(\sum_{\ell\leq L}\max\left\{{Um_{\ell}-T_{\varepsilon}},0\right\}+\frac{dT_{\varepsilon}}{\varepsilon}\right).$ Note that $T_{\varepsilon}$ is non-decreasing with $\varepsilon$. The following lemma then holds.
@@
Now, consider $0<\varepsilon_{1}<\varepsilon_{2}<2d/L$; assume for simplicity that we have $m_{1}\geq m_{2}\geq\ldots\geq m_{L}$. Consider the function $g(\varepsilon,t):=\frac{1}{\sum_{\ell^{\prime}\leq L}m_{\ell^{\prime}}}\cdot\left(\sum_{\ell\leq L}\max\left\{{Um_{\ell}-t},0\right\}+\frac{dt}{\varepsilon}\right).$ Then,
@@ IV-B Performance on Synthetic Datasets
We compare the performances of the following three mechanisms on i.i.d. synthetic datasets: (i) the vanilla Laplace mechanism that releases $M_{\text{Lap}}(\mathcal{D})=f(\mathcal{D})+\mathbf{Z}_{1}$, where $\mathbf{Z}_{1}\sim\text{Lap}\left(\frac{\Delta_{f}}{\varepsilon}\right)$, (ii) the “OPT-worst-case” mechanism that releases $M_{\text{OPT-wc}}(\mathcal{D})=f^{\star}(\mathcal{D})+\mathbf{Z}_{2}$, where $\mathbf{Z}_{2}\sim\text{Lap}\left(\frac{\Delta_{f^{\star}}}{\varepsilon}\right)$, and (iii) the “AKMV” mechanism [10, Sec. 3] that releases $M_{\text{AKMV}}(\mathcal{D})=\widehat{f}_{T}(\mathcal{D})+\mathbf{Z}_{3}$, where $\mathbf{Z}_{3}\sim\text{Lap}\left(\frac{2\Delta_{\widehat{g}_{T}}}{\varepsilon}\right)$, and $\widehat{f}_{T}(\mathcal{D})$ is obtained by clipping the sum of samples of each user to lie in $[0,T]$, where $T$ is the estimate of the $\left\lceil\left(\frac{2}{\varepsilon}\right)\right\rceil^{\text{th}}$-largest sample among $\{\sigma^{(\ell)}\}$, which is estimated with privacy budget $\varepsilon/2$.
The“ average-case” errors on i.i.d. datasets are defined as $\overline{E}_{\text{Lap}}=\mathbb{E}[|\mathbf{Z}_{1}|],\ \overline{E}_{\text{OPT-wc}}=\mathbb{E}[|M_{\text{OPT-wc}}(\mathcal{D})-f(\mathcal{D})|],$ and $\overline{E}_{\text{AKMV}}=\mathbb{E}[|M_{\text{AKMV}}(\mathcal{D})-f(\mathcal{D})|],$ where the expectations are over the randomness in the data samples and in the DP mechanism employed. These errors are then estimated via Monte-Carlo averaging over $10^{4}$ iterations. As a pre-processing step, we replace each of the samples $\{x_{j}^{(\ell)}:\ j\in[m_{\ell}]\}$ of every user $\ell\in[L]$ by the sample average $\frac{1}{m_{\ell}}\cdot\sum_{j\leq m_{\ell}}x_{j}^{(\ell)}$, so as to allow for improved performance post clipping.

### v3

@@ III-A On Clipping Strategies
We work with estimators $\overline{f}=\overline{f}_{\left\{a_{j}^{(\ell)},b_{j}^{(\ell)}\right\}}$ of $f$ obtained by bounding user contributions as follows: for each $\ell\in[L]$ and $j\in[m_{\ell}]$, we let $\overline{\mathbf{x}}_{j}^{(\ell)}:={\mathsf{A}_{a_{j}^{(\ell)},b_{j}^{(\ell)}}}\left(\mathbf{x}_{j}^{(\ell)}\right)$, for reals $0\leq a_{j}^{(\ell)}\leq b_{j}^{(\ell)}\leq U$. In words, $\overline{\mathbf{x}}_{j}^{(\ell)}$ is a projection of $\mathbf{x}_{j}^{(\ell)}$ onto the set $\mathsf{A}_{a_{j}^{(\ell)},b_{j}^{(\ell)}}$, which, intuitively, reduces the range of values that $\mathbf{x}_{j}^{(\ell)}$ can take, and hence its sensitivity too. Figure 1 shows a pictorial depiction of the annulus $\mathsf{A}_{a_{j}^{(\ell)},b_{j}^{(\ell)}}$ when $d=2$ and examples of projections onto the annulus. We then set
$\overline{f}=\overline{f}(\mathcal{D}):=\frac{1}{\sum_{\ell^{\prime}=1}^{L}m_{\ell^{\prime}}}\cdot\sum_{\ell=1}^{L}\sum_{j=1}^{m_{\ell}}\mathbf{\overline{x}}_{j}^{(\ell)}.$
@@ III-B On Worst-Case Errors of Clipping Strategies
For a sample $\mathbf{x}_{j}^{(\ell)}\in\mathsf{A}_{a_{j}^{(\ell)},b_{j}^{(\ell)}}$, it is clear that $\mathsf{A}_{a_{j}^{(\ell)},b_{j}^{(\ell)}}(\mathbf{x}_{j}^{(\ell)})=\mathbf{x}_{j}^{(\ell)}$. Now, suppose that $\mathbf{x}_{j}^{(\ell)}\in\Delta_{U}\setminus\Delta_{b_{j}^{(\ell)}}$. Following Lemma III.1, if $\mathbf{y}=\mathsf{A}_{a_{j}^{(\ell)},b_{j}^{(\ell)}}(\mathbf{a})$, we must have
$\lVert\mathbf{y}-\mathbf{x}_{j}^{(\ell)}\rVert_{1}=\sum_{i\leq d}((x_{j}^{(\ell)})_{i}-y_{i})=\lVert\mathbf{x}_{j}^{(\ell)}\rVert_{1}-b_{j}^{(\ell)}.$
Hence, the worst-case clipping error for sample $\mathbf{x}_{j}^{(\ell)}$ is
$\max\ \lVert\mathsf{A}_{a_{j}^{(\ell)},b_{j}^{(\ell)}}(\mathbf{x}_{j}^{(\ell)})-\mathbf{x}_{j}^{(\ell)}\rVert_{1}=U-b_{j}^{(\ell)}.$
By symmetric arguments, one can show that if $\mathbf{x}_{j}^{(\ell)}\in\Delta_{a_{j}^{(\ell)}}\setminus\delta_{a_{j}^{(\ell)}}$, we must have
$\max\ \lVert\mathsf{A}_{a_{j}^{(\ell)},b_{j}^{(\ell)}}(\mathbf{x}_{j}^{(\ell)})-\mathbf{x}_{j}^{(\ell)}\rVert_{1}=a_{j}^{(\ell)}.$
Thus, overall, we obtain that
@@
Via arguments entirely analogous to that in the proof of Lemma III.2, the user-level sensitivity of $\overline{f}$ is
$\Delta_{\overline{f}}=\frac{\max_{\ell\leq L}\sum_{j\leq m_{\ell}}\left(b_{j}^{(\ell)}-a_{j}^{(\ell)}\right)}{\sum_{\ell^{\prime}\leq L}m_{\ell^{\prime}}},$
since in the worst-case, all samples of a user $\ell$ are changed each from $(a_{j}^{(\ell)},0,\ldots,0)$ to $(b_{j}^{(\ell)},0,\ldots,0)$. Using $\mathbb{E}[\lVert\mathbf{Z}\rVert_{1}]=d\mathbb{E}[|Z_{1}|]=d\Delta_{\overline{f}}/\varepsilon$ gives us the required result. ∎
@@
The optimization problem in (5) is a linear programming problem. By standard arguments via the necessity of the KKT conditions [18, Sec. 5.5.3], there must exist reals $\lambda_{\tau}\geq 0$ and $\lambda_{\ell},\ \mu_{\ell}\geq 0$, for each $\ell\in[L]$ (or Lagrange multipliers), such that the function
($\mathcal{L}(\{S^{(\ell)}\},$ $\lambda_{S},\{\lambda_{\ell},\mu_{\ell}\}):=\sum_{\ell\leq L}S^{(\ell)}+\frac{S}{\varepsilon}-\lambda_{S}S\ +$
$\sum_{\ell}\lambda_{\ell}\cdot\left(Um_{\ell}-2S^{(\ell)}-S\right)-\sum_{\ell}\mu_{\ell}S^{(\ell)}$
obeys the following properties.
@@
  • Complementary slackness: We have that
    $\lambda_{\tau}\tau=0,\ \lambda_{\ell}\cdot\left(Um_{\ell}-2S^{(\ell)}-\tau\right)=0,\ \text{and}\ \mu_{\ell}S^{(\ell)}=0,$
    for all $\ell\in[L]$.
@@
Observe that from Theorem III.1 and (4), the optimal worst-case error is
$E^{\text{OPT}}(\varepsilon)=\frac{1}{\sum_{\ell^{\prime}\leq L}m_{\ell^{\prime}}}\cdot\left(\sum_{\ell\leq L}\max\left\{\frac{Um_{\ell}-T_{\varepsilon}}{2},0\right\}+\frac{dT_{\varepsilon}}{\varepsilon}\right).$
Note that $T_{\varepsilon}$ is non-decreasing with $\varepsilon$. The following lemma then holds.
@@
Now, consider $0<\varepsilon_{1}<\varepsilon_{2}<2d/L$; assume for simplicity that we have $m_{1}\geq m_{2}\geq\ldots\geq m_{L}$. Consider the function $g(\varepsilon,t):=\frac{1}{\sum_{\ell^{\prime}\leq L}m_{\ell^{\prime}}}\cdot\left(\sum_{\ell\leq L}\max\left\{\frac{Um_{\ell}-t}{2},0\right\}+\frac{dt}{\varepsilon}\right).$ Then,
@@ IV-B Performance on Synthetic Datasets
We compare the performances of the following three mechanisms on i.i.d. synthetic datasets:
  1. 1.
    the vanilla Laplace mechanism that releases $M_{\text{Lap}}(\mathcal{D})=f(\mathcal{D})+\mathbf{Z}_{1}$, where $\mathbf{Z}_{1}\sim\text{Lap}\left(\frac{\Delta_{f}}{\varepsilon}\right)$;
  2. 2.
    the “OPT-worst-case” mechanism that releases $M_{\text{OPT-wc}}(\mathcal{D})=f^{\star}(\mathcal{D})+\mathbf{Z}_{2}$, where $\mathbf{Z}_{2}\sim\text{Lap}\left(\frac{\Delta_{f^{\star}}}{\varepsilon}\right)$; and
  3. 3.
    the “AKMV” mechanism [10, Sec. 3] that releases $M_{\text{AKMV}}(\mathcal{D})=\widehat{f}_{T}(\mathcal{D})+\mathbf{Z}_{3}$, where $\mathbf{Z}_{3}\sim\text{Lap}\left(\frac{2\Delta_{\widehat{g}_{T}}}{\varepsilon}\right)$, and $\widehat{f}_{T}(\mathcal{D})$ is obtained by clipping the sum of samples of each user to lie in $[0,T]$. Here, $T$ is the estimate of the $\left\lceil\left(\frac{2}{\varepsilon}\right)\right\rceil^{\text{th}}$-largest sample among $\{\sigma^{(\ell)}\}$, which is estimated with privacy budget $\varepsilon/2$.
The“ average-case” errors on i.i.d. datasets are defined as
$\overline{E}_{\text{Lap}}=\mathbb{E}[|\mathbf{Z}_{1}|],\ \overline{E}_{\text{OPT-wc}}=\mathbb{E}[|M_{\text{OPT-wc}}(\mathcal{D})-f(\mathcal{D})|],$
and
$\overline{E}_{\text{AKMV}}=\mathbb{E}[|M_{\text{AKMV}}(\mathcal{D})-f(\mathcal{D})|],$
where the expectations are over the randomness in the data samples and in the DP mechanism employed. These errors are then estimated via Monte-Carlo averaging over $10^{4}$ iterations. As a pre-processing step, we replace each of the samples $\{x_{j}^{(\ell)}:\ j\in[m_{\ell}]\}$ of every user $\ell\in[L]$ by the sample average $\frac{1}{m_{\ell}}\cdot\sum_{j\leq m_{\ell}}x_{j}^{(\ell)}$, so as to allow for improved performance post clipping.